Calculate the MCV, MCH, and MCHC indices given the following information.

1) RBC: 5.0 X 1012 HCT: 45% HGB: 16.0 g/dL

2) RBC: 2.3 X 1012 HCT: 22% HGB: 7.3 g/dL

3) RBC: 6.1 X 1012 HCT: 50% HGB: 18 g/dL
Answer #1

a) MCV (Mean cell volume) : MCV = (Haemocrit %/ 100)/RBC(1012 l) Normal range : 82-98 fL

1) MCV = 45 % / 100)/5.0 x 1012 = 0.09 x 10-12 = 90 x 10-15 l = 90 fL

2) MCV = 22 % / 100)/2.3 x 1012 = 0.095 x 10-12 = 95 x 10-15 l = 95 fL

3) MCV = 50 % / 100)/6.1 x 1012 = 0.082 x 10-12 = 82 x 10-15 l = 82 fL

b) MCH (Mean cell haemoglobin) : (HGB x 10)/RBC(1012 l) Normal range : 27 -33

1) MCH = (16 g/dL x 10)/5.0 x 1012 = 32 x 10-12 g = 32 pg (picogram)

2) MCH = (7.3 g/dL x 10)/2.3 x 1012 = 31.7 x 10-12 g = 31.7 pg

3) MCH = (18 g/dL x 10)/6.1 x 1012 = 29.5 x 10-12 = 29.5 pg

c) MCHC (Mean cell haemoglobin concentration) : MCHC = (HGB (g/dL) x 100)/(haematocrit % ) Normal range : 30-35 g/dl

1) MCHC = (16 g/dl x 100)/45% = 35.55 g/dl

2) MCHC = (7.3 g/dl x 100)/22 % = 33.18 g/dl

3) MCHC = (18 g/dl x 100)/50% = 36 g/dl